# Wave-induced drift in third-order deep–water theory

### v1

@@ 4 Bichromatic waves
In the previous section we considered monochromatic waves. The term “monochromatic” refers to the fact that a single mode is assumed in the lowest order problem. As we have seen, higher harmonics are excited at each order of the perturbation expansion, so that the spectrum does not, in fact, remain monochromatic. Far from being a purely mathematical phenomenon, this harmonic generation is observed every time a wavemaker is excited in a flume, and remains an active area of experimental research (Khait & Shemer, 2019). Consequently, when we speak of bichromatic waves, this means that two modes and are assumed in the lowest order problem. The second order solution then contains sum and difference terms of any two of these, while the third order contains sum and difference terms of any three, with an attendant increase in complexity. Up to second order the solution is contained in the work of Dalzell (1999), while the third order solution for bichromatic waves appears first in the work of Madsen & Fuhman (2006). We make use of the equivalent formulation of Section 2.3.
4.1 Linear theory
We start with the simplest scenario, where we retain only linear terms in the bichromatic velocity potential
To understand what this means for particle kinematics, we are again confronted with two options: solve the particle trajectory mapping numerically or approximate it in the same fashion as the monochromatic Stokes’ drift in Section 3.
Exactly as for monochromatic waves, if the lowest order terms in a Taylor series expansion of (37) are retained, the resulting system with explicit time dependence
is trivially integrated. Substituting the solution into the Taylor series expansion (cf. (29)) leads to a somewhat lengthy term; while the products of the first order solution and the velocity field produce terms involving the time dependence of the equations is, of course, explicit. Upon integration the secular terms lead immediately to the drift velocity
which is simply the sum of the Stokes’ drifts for modes and over a time t. This simple idea, that the Stokes drift of a linear sum of waves is the sum of the Stokes drifts of the waves underlies the successful application of Stokes drift to spectra as pioneered by Kenyon (1969).
As we have for monochromatic waves, it is worth comparing the Stokes drift formulation (40)with the actual particle trajectory mapping. Such a comparison is shown in Figure 5, for different snapshots of one bichromatic wave-train. Although it is mathematically elementary, it is worth some digression to explain this figure. Bichromatic wave-trains look quite different in space and in time, so we choose to visualise the free surface as a function of time at different spatial locations. Two radian frequencies rad/s and rad/s are considered, such that their superposition has a (Eulerian, temporal) period of s. (With m/s2 the bichromatic wave train also has a spatial period of approximately 986.2 m.) Several snapshots of the free-surface time-series over one such period are shown in Figure 5 panels (b)–(d), with the envelope indicated by the pair of dashed, black curves.
Panel (a) of Figure 5 shows the trajectories of particles starting at and various depths ( and -4.8 m) over one temporal period. Exactly as in Figure 2, the solid curves are obtained by integrating the particle trajectory mapping with the bichromatic velocity field obtained from (37), while the dashed curves shown an analytical solution obtained from retaining the 2nd order terms in the Taylor series expansion of the same equations. Consequently, the forward movement of a particle along the dashed trajectory is given simply by (40), shown as a black, dashed line in Panel (a), while the forward movement of a particle computed from numerical integration is shown as a black solid line
Panels (b)–(d) show the free surface elevation at three spatial locations, and the panels below each show the corresponding trajectories (panel (e) shows an enlarged version of the blue trajectories in panel (a) with initial position ). Note that, while the free surface envelope in panels (b) and (d) may look very similar, the trajectory in (e) begins at a crest while the trajectory in (g) begins near a trough. This difference is clearly visible in the trajectories themselves. Trajectory (f), on the other hand, begins near a local minimum (in space) close to the mean water level
As for monochromatic waves, we can compare the computed Stokes drift (40) with the values obtained directly from integration of the particle trajectory ODEs. For the example considered in Figure 5 we observe (as in Section 3) that the Lagrangian period is always larger than the Eulerian period. Averaging the drift velocity over one spatial period yields Figure 6, where dashed curves represent the Stokes drift velocity and solid curves the computed averaged Lagrangian drift velocity. Bichromatic wave-trains with rad/s and rad/s and and 0.1 are considered. The same qualitative features as in Figure 3 are observed: linear theory predicts a larger forward drift than the Stokes’ drift approximation. This is also the case when the bichromatic wave modes have different slopes.
4.2 Nonlinear theory
4.2.1 Second order
If linear bichromatic waves are very simple, the second order expression obtained from (22) is not much more complicated:
@@ 4.2.1 Second order
where and the sum harmonic vanishes for unidirectional deep water waves (the same is not true of the free surface elevation, whose prominent sum harmonics can be seen in Figure 1).
@@
The second order solution (41) lends itself to a calculation of the Stokes drift as outlined in Section 3.1: calculate , Taylor expand about a point , keep only the constant terms in the Taylor expansion and calculate approximate trajectories , subsequently substitute these trajectories into the linear terms in the Taylor expansion and integrate the resulting particle trajectory equations in time. The result of this procedure is
which consists of the term (40) from the linear theory as well as a new term (denoted (II)) that arises from the difference harmonics (recall that we assume ).
At first sight the term (II) may seem insignificantly small, however there are regimes where it becomes important in a description of particle drift. It is simplest to analyse this by fixing 1/m and varying for At the top of the fluid domain we observe that the contribution (II) grows monotonically with depicted as the dashed, yellow curve in Figure 7(a). A consequence of this is that there is minimum of the total 2nd order Stokes drift (42) for a particular value of For we find that term (II) has a single maximum and subsequently decays, as shown for two depths in Figure 7(b)–(c). As we descend into the fluid, the maximum of term (II) moves towards smaller values of
@@
While overall particle motion decreases the deeper we descend into the fluid, there are regions where the 2nd order contribution to Stokes’ drift has a sizable effect (see Figure 7(b)) and in fact dominates the 1st order contribution (I), as shown around in Figure 7(c). This is in line with the importance of difference-harmonic terms in generating dynamic pressures that reach the ocean bed (Longuet-Higgins, 1950).
To further understand the importance of the second order Stokes’ drift correction, we compare the Stokes’ drift formulation with the particle drift obtained from integration of the particle trajectory mapping. Panels (d)–(f) of Figure 7 show the average horizontal drift obtained by integrating using the 1st order velocity field (37) or the 2nd order velocity field (41). The averaging is performed over initial conditions spanning one wavelength of the bichromatic envelope, and both modes have identical steepness .
Panel (d) depicts the drift at for a bichromatic wave train with 1/m and 1/m. The discrepancy between the linear Stokes drift (I) and 1st order solution has already been noted in Figure 6, and is seen again in comparing the solid and dashed blue lines. It is therefore rather surprising that incorporating the second order Stokes drift (42) leads to much better agreement with the exact, numerical solution, as shown by the dashed and solid red curves.
In panels (e) and (f) of Figure 7 we consider a bichromatic wave train with 1/m and 1/m, close to the maximum of the yellow curve (II) shown in panels (b) and (c) at depths of m and -5 m, respectively. At such depths the overall particle motion is much smaller, but for a range of wavenumbers the 2nd order component dominates the Stokes’ drift. Integrating the particle trajectory mapping bears this out: the blue curves showing the expression (I) overlap with the solutions of the particle trajectory mapping with 1st order velocity field, as in Figure 6. Significantly, the red lines showing the analytical expression (I)+(II) given in (42) also overlap with the respective 2nd order solutions. It can be checked that for where contribution (II) is insignificant, all four lines overlap.
4.2.2 Difference harmonics and particle drift
@@
Why do we observe this reversal, where linear theory predicts a larger drift near the surface, but second order theory predicts a larger drift at depth? The answer lies in the structure of the difference harmonic terms, and as a prelude to our work on wave groups it is expedient to investigate their role in bichromatic seas.
The first consideration should be of the velocity field which drives the particle trajectory mapping. We consider the waves studied in Figure 7 with and 1/m, and calculate horizontal velocities directly from (41). When only the first order contributions are retained we find the familiar pattern of forward velocity below the crests and backward velocity below the troughs, with an exponential fall-off in depth as depicted in the top panel of Figure 8.
@@
If instead we focus only on the second-order contributions , the situation looks quite different. The exponential decay in velocities is slower because is larger than either or for and we find a negative horizontal velocity under the centre of the bichromatic wave train, as shown in the bottom panel of Figure 8. The full second-order picture is a sum of the first and second order terms, shown in the middle panel of the same figure, and it is clear that the first order dominates near the surface while the second-order difference harmonic dominates at depth. This can be observed clearly by considering the zero contour lines of the horizontal velocity, shown in black for each case.
@@
The generic situation for particle paths in a bichromatic wave train is thus as follows: close to the surface the first order contributions dominate, giving the curlicue particle paths observed in Figure 5 and in Figure 9(b). The effect of difference harmonics is to retard slightly the forward motion of the particles, due to the slow moving (in deep water the group velocity is half the phase velocity) region of backward (negative) velocity under the highest crests observed in Figure 8. This retardation from second-order difference terms gives rise to the larger drift in the linear theory, observed in Figure 9(c).
@@
Deeper into the fluid column the effect of the first order free-modes becomes negligible. The fluid motion is dominated by the difference harmonic term , and the particle paths behave essentially as those of a monochromatic wave with that wavenumber. This is clearly illustrated in Figure 9(d), where we see that the modes and which make up the first-order theory are only a small perturbation on the displacement due to the difference harmonic term.
@@
It is worth pointing out the existence of a region beneath the bichromatic wave train where the particle displacement is largely against the direction of wave propagation, as seen between in Figure 9(e), and indicated with yellow circles. This should not be surprising, since it simply reflects the effective motion of the difference harmonic term as shown in panel (d). Despite the existence of this ‘return flow’ – which we shall encounter again in our discussion of wave groups – the net particle motion remains forward at every depth in the fluid, as shown in Figure 7 and Figure 10 below. Indeed, the relatively large displacement associated with difference harmonic terms at depth gives rise to the larger drift observed for the second-order solution.
4.2.3 Third order
A confluence of the effects of dispersion correction and harmonic generation occurs for the first time at third order. For the frequency correction appearing there we have the notably asymmetric equations
for which were first found by Longuet-Higgins & Phillips (1962). Note the presence of the Stokes correction found in (35), and the fact that the longer wave () has a much more significant effect on the shorter wave () than vice versa.
The higher harmonics (or bound waves) have the well-known effect of sharpening wave crests and flattening wave troughs, as shown in Figure 1. To first order (dark blue curve) this figure shows the same bichromatic wave-train treated in Section 4.1, albeit with the slope of each mode increased to for ease of visualisation. Note that the 1st and 2nd order solutions (shown in dark and light blue) employ linear dispersion. The 3rd order solution (shown in red) includes the mutual frequency correction (43)–(44), and so has a different period than – and in fact is seen to move out of phase with – the lower order solutions.
While a compact form of the third-order bichromatic potential akin to (41) is not available, it is relatively simple to simulate the exact, third-order theory on a computer using the results of Section 2.3. This allows us to compare the computed forward drift for the same bichromatic wave train using the 1st order velocity field, the 2nd order velocity field, or the 3rd order velocity field, as well as the various formulations of Stokes drift. Such a comparison is shown in Figure 10, which considers the depth-dependence of the wave trains shown in Figure 7. Because the particle drift depends on the initial horizontal position – i.e. at which point in the phase the wave train is encountered – this figure is calculated by averaging over one spatial period at each depth.
Close to the surface we observe the largest drift in 1st order theory, followed by 2nd order and 3rd order. Incorporating difference harmonic terms (II) into the Stokes drift leads to a closer agreement with the exact second and third order drift as shown in the left panel of Figure 10. With greater depth, and for nearby wavenumbers, we have a distinct reversal in the ordering: 1st order theory, whether exact or using the Stokes drift approximation, yields the smallest forward drift. The 2nd order theory gives the largest drift, with the 3rd order slightly smaller. Once again, incorporating difference harmonic terms into the Stokes drift provides a distinct improvement in such cases over depth.
@@ 5.1 Focused linear wave groups
In treating waves with an increasing number of linear harmonic constituents we also have a plethora of parameters that can be tuned. One case of interest, from both an experimental and theoretical standpoint is that of a (focused) wave group. Linear focusing is quite straightforward: if our free surface is a superposition of sinusoids as per (24), then writing
@@
readily gives a wave train that focuses – i.e. all components come into phase – at . Pulse-like wave groups can be created by suitable tuning of the amplitude spectrum One possible choice is a Gaussian amplitude spectrum of the form
@@
An example of a wave train with such an amplitude spectrum is shown in Figure 11(a), where we have chosen equidistant modes between and 1/m, , and The focusing location is chosen to be , and the focus occurs at a crest.
In order to more clearly visualise the behaviour of the particles the curves in Figure 11 have been coloured with a time-varying gradient. Comparing panels (a) and (b) we observe that a particle at at time initially undergoes very little motion. Panel (c) shows that its horizontal position is effectively constant at until nearly s, as is its vertical position (not shown). As the group passes the particle moves with the wave motion, before settling at a new horizontal position close to m.
At greater depth the particle motion from the linear wave group is shown in Figure 11(d)–(e) where m. Here the particle seems very nearly to return to its initial position along a spiraling trajectory shown in panel (d), with nearly vanishing forward drift, as expected from our discussion of monochromatic and bichromatic waves.
5.2 Nonlinear focused waves
The particle trajectories obtained by integrating the linear velocity field in Figure 11 can easily be recalculated using either second or third-order theory. Experience with bichromatic waves in Section 4 teaches us to expect that all theories will behave similarly near the free surface, with linear theory giving the largest forward drift followed by second and third order. Indeed this is what we observe, as shown in Figure 12(a)–(b), where the colour gradient is the same as in Figure 11.
In contrast to the bichromatic case, we now have a plethora of difference harmonic terms which have an influence further down the water column. Difference harmonics close to for the group dominate, and lead to a 25-fold increase in the displacement at depth when compared to the linear theory, as shown in Figure 12(c)–(d). We also observe the return flow associated with these trajectories, exactly as for bichromatic waves in Section 4.2.2. Interestingly, the third-order contribution, consisting of both a change in dispersion relation as well as the presence of additional harmonics, has a notable effect on the particle displacement throughout.
We can get a better idea of the depth-dependence of the particle motion from Figure 13, which shows the horizontal position of a particle initially at m after the passage of the wave group discussed above. First, second and third-order solutions are denoted by blue, red and yellow curves, respectively, and the Stokes drift (generalising (40) to 10 modes) is denoted by a purple curve with label ’S’. This figure clearly illustrates the role of higher-order theories in reducing the drift close to the surface, and in inducing a significant return flow beneath the group with depth. Indeed, while linear theory and Stokes drift are reasonably close in Figure 13(b), they cannot capture the sizable negative velocities beneath the group.
Care should be taken, however, not to overinterpret this result. As we have seen for bichromatic waves in Section 4.2.2, the return flow is a highly localised phenomenon, and does not indicate a net negative drift. Our wave groups are clearly not envelope solitary waves – whose existence in the third-order water wave problem remains open (Slunyaev & Shrira, 2013) – composed as they are of a finite number of Fourier modes. Consequently there is no approach to a limiting state of zero-flow as the wave group passes, rather the passage of one group simply heralds the appearance of another following behind it. The particle motion induced by such a succession of wave groups recovers the expected forward drift.
5.3 Random waves

### v2

@@ 4 Bichromatic waves
The simplest “irregular” waves, corresponding to unsteady flows, are those that consist of two harmonics and at first order, and are therefore sometimes termed “bichromatic” waves. The second order solution then contains sum and difference terms of any two of these, while the third order contains sum and difference terms of any three, with an attendant increase in complexity. Up to second order the solution is contained in the work of Dalzell (1999), while the third order solution for bichromatic waves appears first in the work of Madsen & Fuhman (2006). We make use of the equivalent formulation of Section 2.3.
Up to second order the velocity potential can be written as the remarkably simple expression
@@
where without loss of generality, and the sum harmonic vanishes for unidirectional deep water waves (the same is not true of the free surface elevation, whose prominent sum harmonics can be seen in Figure 1).
We show the free-surface time series of one such bichromatic solution in Figure 5(a), where the inclusion of 2nd and 3rd order terms can be seen to have only a minor effect. This is not the case when considering the particle paths, which exhibit notable differences between linear and higher-order theories, particularly at depth. Figure 5(b) shows the paths of particles close to the surface with initial position (marked by ) and different final positions (marked with , and in the respective colours).
Note that Figures 5(b),(d) are simply obtained by integrating where denotes the velocity at a given order ( or 3). The relative importance of each term can be assessed by integrating separately for each , which yields Figures 5(c),(e). From these it is clear that the principal contribution at the surface comes from the 1st order velocities, with the principal contribution at depth coming from 2nd order velocities. The 3rd order contributions (in yellow) are barely visible near in panel (e).
@@
The second order solution (36) lends itself to a calculation of a modified Stokes drift in a rather straightforward manner: calculate from (36), Taylor expand about a point , keep only the constant terms in the Taylor expansion and calculate approximate trajectories , subsequently substitute these trajectories into the linear terms in the Taylor expansion and integrate the resulting particle trajectory equations in time. The result of this procedure is
which consists of the sum of the Stokes drifts of each mode separately (I) as well as a new term222While this paper was under review, an article by Liao & Zou (2025) appeared which also derives this term and the extension (41), see (C11) in their Appendix C. (denoted (II)) that arises from the difference harmonics (recall that we assume ). Just as the Stokes drift derived from linear theory includes quadratic terms, so the modified Stokes drift derived from second order theory now includes (formally) quartic terms.
@@
While overall particle motion decreases as we descend into the fluid, there are regions where the 2nd order contribution to Stokes’ drift has a sizable effect (see Figure 6(b)) and in fact dominates the 1st order contribution (I), as shown around in Figure 6(c).
Figure 6 shows the terms (I) and (II) of (37) at three depths: (a) , (b) and (c) for a fixed wave 1/m. It is simplest to analyse the relative importance of the terms by setting and varying We note that the contribution (II) at the top of the fluid domain grows monotonically with , as can be observed in Figure 6(a) (note that the steepness of both modes is fixed at ). We also note that there is a minimum of the total 2nd order Stokes drift at the surface. By contrast, for the quartic term (II) has a single maximum and subsequently decays, as shown in Figure 6(b)–(c). As we descend into the fluid, this maximum moves towards smaller values of
The principal difficulty we encounter – for the first time with bichromatic waves – is connected to the unsteadiness of the flow field and the lack of a unique solution towards which different expansions are known to converge. Unlike the monochromatic waves considered in Section 3, we cannot simply compare waves of the same length and height, which differ geometrically only in the curvature of the profile between crest and trough. The problem is compounded at third order by the appearance of an asymmetric frequency correction (for ) first found by Longuet-Higgins & Phillips (1962)
This means that bichromatic waves at third-order (and 5th, 7th, and so on) will be out of phase with their lower order counterparts, as seen in Figure 1.
Compounded by a paucity of results (theoretical or experimental) on drift associated with bichromatic waves, this makes comparisons difficult. However, we can employ the methodology tested for monochromatic waves in Section 4 to obtain results for bichromatic waves at the surface and at depth, and compare these to the approximate Stokes drift derived in (37).
To obtain drift at the surface we integrate the equation for the horizontal particle position where the velocity field is given by Taylor expansion about to a given order. The 2nd order horizontal velocity can immediately be read off from (36), while the 3rd order velocity, for which a simple algebraic expression has not been found, is obtained from the Hamiltonian expansion of Section 2.
Figure 7 compares the 2nd and 3rd order theory with the classical Stokes drift (SD I) and the modified Stokes drift (SD I+II). In panel (a), where the two waves are close in frequency, we expect the contributions of the terms (II) to be negligible, and indeed the two formulations largely coincide. For more widely separated waves and rad/s, shown in panel (b) the difference harmonic contributions of (II) are more prominent. In both cases the 2nd and 3rd order theories give an enhanced drift, which is consistent with the results obtained for monochromatic waves and shown in Figure 3.
We have also compared with the 4th order drift obtained from the 3rd order bichromatic solution in Lagrangian variables obtained in Appendix B of Blaser et al. (2025), which generalises Pierson’s (1961) classical 2nd order result. Unfortunately, bichromatic waves in Eulerian and Lagrangian coordinates are not the same; rather, the wave profiles are geometrically distinct, with a phase-shift that occurs between successive crests of both solutions, as already noted by Fouques & Stansberg (2008)333This phase shift is not due to differences in dispersion relation, as might be hypothesised from Pierson (1961), since Blaser et al (2025) recover the Eulerian dispersion correction (38)–(39) at 3rd order.. Whether it is possible to adapt the Lagrangian perturbation expansion to match the Eulerian solution (or vice versa) remains an open question. For well-separated frequencies (panel (b)) the Eulerian and Lagrangian solution remain remarkably close, while for two nearby frequencies (panel (a)) the notable difference grows with wave slope – and is visible also when comparing the free surface plots of the two solutions.
Below the surface, exactly as shown for monochromatic waves, it is possible to use the particle trajectories obtained by integrating the Eulerian velocity field to obtain the Stokes drift. This is put to the test in Figure 8, where we select a bichromatic wave train with significant subharmonic drift as per (37), namely and 1/m. This corresponds to the circles shown in Figure 6.
Figure 8 compares 1st, 2nd and 3rd order theory with the formulations (I) and (I)+(II) of (37), as well as the 4th order Lagrangian drift of Blaser et al. (2025). In fact, the particle paths obtained from integrating 1st order theory coincide completely with the classical Stokes drift (SD I). On the other hand, there is very good agreement between 2nd and 3rd order theories, the improved Stokes drift formulation (SD I+II) and the fourth order Lagrangian theory (L4). The same comparison for a pair of more widely separated wavenumbers (such as and 1/m, not shown) would demonstrate that all six curves essentially coincide, as might be predicted from Figure 6 (noting that there are no notable flows induced by the difference harmonic).
4.1 Difference harmonics and particle drift
@@
In general, we see the importance of leading order effects near the surface of bichromatic waves, while second order effects become more prominent at depth. This is no surprise, and follows from the structure of the difference harmonic terms themselves. As a prelude to our work on wave groups it is expedient to investigate their role in bichromatic seas.
The first consideration should be of the velocity field which drives the particle trajectory mapping. We consider the waves studied in Figure 6 with and 1/m, and calculate horizontal velocities directly from (36). Indeed, Figure 5 makes clear that the 3rd order velocity field – though difficult to calculate – makes an essentially negligible contribution. In fact, the only material contributions of the 3rd order theory are due to the dispersion correction (38)–(39), as can be seen by comparing panels (b) and (d) with panels (c) and (e) in Figure 5. These dispersion corrections do not affect the instantaneous velocity field, only its time evolution.
@@
When only the first order contributions are retained we find the familiar pattern of forward velocity below the crests and backward velocity below the troughs, with an exponential fall-off in depth as depicted in the top panel of Figure 9. If instead we focus only on the second-order contributions , the situation looks quite different. The exponential decay in velocities is slower because is larger than either or for and we find a negative horizontal velocity under the centre of the bichromatic wave train, as shown in the bottom panel of Figure 9. The full second-order picture is a sum of the first and second order terms, shown in the middle panel of the same figure, and it is clear that the first order dominates near the surface while the second-order difference harmonic dominates at depth. This can be observed clearly by considering the zero contour lines of the horizontal velocity, shown in black for each case.
@@
The generic situation for particle paths in a bichromatic wave train is thus as follows: close to the surface the first order contributions dominate, giving the curlicue particle paths observed in Figure 5(b). The effect of difference harmonics is to retard slightly the forward motion of the particles, due to the slow moving (in deep water the group velocity is half the phase velocity) region of backward (negative) velocity under the highest crests observed in Figure 5(e).
@@
Deeper into the fluid column the effect of the first order free-modes becomes negligible. The fluid motion is dominated by the difference harmonic term , and the particle paths behave essentially as those of a monochromatic wave with that wavenumber. This is clearly illustrated in Figure 5(e), where we see that the modes and which make up the first-order theory are only a small perturbation on the displacement due to the difference harmonic term.
@@
It is worth pointing out the existence of a region beneath the bichromatic wave train where the particle displacement is largely against the direction of wave propagation, as seen between in Figure 5(e) (see markers ’+’). This should not be surprising, since it simply reflects the effective motion of the difference harmonic term as shown in panel (d). Despite the existence of this ‘return flow’ – which we shall encounter again in our discussion of wave groups – the net particle motion remains forward at every depth in the fluid, as shown in Figure 6 and Figure 8 below. Because the solution is periodic, over many periods the net effect of difference harmonics is to enhance forward drift, as shown in Figure 8.
@@ 5.1 Focused linear wave groups
In treating waves with an increasing number of linear harmonic constituents we also have a plethora of parameters that can be tuned. One case of interest, from both an experimental and a theoretical standpoint is that of a (focused) wave group. Linear focusing is quite straightforward: if our free surface is a superposition of sinusoids as per (25), then writing
@@
readily gives a wave train that focuses – i.e. all components come into phase – at . Pulse-like wave groups can be created by suitable tuning of the amplitude spectrum
One possible choice is a Gaussian amplitude spectrum of the form
@@
An example of a wave train with such an amplitude spectrum is shown in Figure 10(a), where we have chosen equidistant modes between and 1/m, , and The focusing location is chosen to be , and the focus occurs at a crest. Following Blaser et al. (2025) we define
as the maximum slope for the linear focused wave, and use this as a measure of nonlinearity going forward.
The particle behaviour in Figure 10(b),(d) is in many ways similar to simpler, bichromatic wave group shown in Figure 5. Towards the top of the fluid column, 1st order theory dominates, while at depth the effect of the in-phase subharmonics associated with higher-order theories clearly accounts for the bulk of the particle motion. In order to further illustrate this point, panels (c) and (e) show only the horizontal component of the particle position as a function of time , as the wave group passes over the initial position. Towards the surface (c) there is nearly no motion until s, while at depth we observe a steady forward drift in the positive -direction. Under the centre of the group, as expected, we see a strong forward drift towards the surface (panel c) and backward (return) flow at depth (panel e).
This effect can be seen even more clearly in Figure 11, which illustrates (for a single particle) the depth dependence of the horizontal motion. The (leading order) free surface of a focused Gaussian wave train with 1/m is shown in panel (a), in both the crest focused (blue, solid curve) and trough focused (red, dash-dotted curve) case. Panel (b) shows the horizontal displacement of a particle as the wave group passes it.
Interestingly, there is a marked difference in the return flow between crest and trough focused cases, with even 1st order theory showing some return flow for trough focused waves (dash-dotted curves). Such (localised) return flows are not captured by Stokes drift approximations, whether based simply on the linear theory or incorporating second order effects via the extension of (37)
Indeed, the effects shown in Figure 11 are highly phase-dependent, and rely on suitable choice of position in the spatio-temporal evolution of the group, while the approximate formulae (41) can only capture an averaged drift.
As we have done for monochromatic and bichromatic waves, we can apply 2nd and 3rd order, deep-water theory directly to calculate the surface drift of focused wave groups. The results of such a computation are reported in Figure 12, which is the counterpart of Figures 3 and 7. In this case, we have initialised our wave group with a central frequency and using modes have defined the amplitudes where the frequencies are chosen via for bandwidth and are the wavenumbers obtained from the linear dispersion relation. This is precisely the formulation used by Blaser et al. (2025), and allows us to compare our results with their theoretical and numerical values.
Following the trends for monochromatic and bichromatic waves, we note that the lowest order Stokes drift formulation (SD I, which treats the Stokes drift as a sum of the drifts of each individual mode (32)) gives the smallest value. The 2nd and 3rd order formulations show good agreement with both laboratory experiments () and fully nonlinear simulations () up to rather high steepness , even though our formulation neglects amplitude evolution which is included in the numerical simulations (see also the discussion in Appendix C).
5.2 Random multichromatic waves